OpenStudy (el_arrow):

find the area of the surface obtained by rotating the curve about the x-axis.

OpenStudy (el_arrow):

x=1/3(y^2+2)^3/2, 1<y<3

OpenStudy (el_arrow):

hello

OpenStudy (perl):

hi

OpenStudy (el_arrow):

could you help me with this problem

OpenStudy (perl):

the form of an integral rotating about x axis is $$\huge\int 2\pi~ y ~ ds$$

OpenStudy (el_arrow):

okay so how do i get y and ds

OpenStudy (perl):

you have two options $$\huge\int 2\pi~ f(x) ~ \sqrt{1+(dy/dx)^2}dx$$ or $$\huge\int 2\pi~ y ~ \sqrt{1+(dx/dy)^2}dy$$

OpenStudy (el_arrow):

can we do the second one?

OpenStudy (perl):

yes

OpenStudy (el_arrow):

or whichever is the easiest

OpenStudy (perl):

that will be easiest since we were given x = g(y)

OpenStudy (el_arrow):

okay

OpenStudy (perl):

so first find dx/dy

OpenStudy (el_arrow):

i got 1/2 * (y^2+2)^1/2 * 2y

OpenStudy (perl):

$$x=1/3(y^2+2)^{3/2} \\ \frac{dx}{dy} = 1/3 *(3/2)(y^2+2)^{1/2} * 2y $$

OpenStudy (perl):

did i copy x(y) correctly?

OpenStudy (el_arrow):

yes that right

OpenStudy (perl):

ok i got same as you

OpenStudy (perl):

$$\huge \int_{1}^{3} 2\pi y \sqrt{ 1+(y*(y^2+2)^{1/2} )^2}~dy $$

OpenStudy (el_arrow):

okay

OpenStudy (el_arrow):

can we multiply the y in

OpenStudy (el_arrow):

okay but what happened to the 1/2?

OpenStudy (perl):

$$ \Large x=1/3(y^2+2)^{3/2} \\ \Large \frac{dx}{dy} = 1/3 *(3/2)(y^2+2)^{1/2} * 2y \\ \Large \frac{dx}{dy}= 1/3 * 3/2 * 2 y (y^2+2)^{1/2} \\ \Large \frac{dx}{dy}= y (y^2+2)^{1/2} $$

OpenStudy (el_arrow):

alright got that

OpenStudy (el_arrow):

so now all we need is the y right?

OpenStudy (perl):

$$ \Large \int_{1}^{3} 2\pi y \sqrt{ 1+(y*(y^2+2)^{1/2} )^2}~dy \\ \Large \int_{1}^{3} 2\pi y \sqrt{ 1+y^2(y^2+2)}~dy \\\Large \int_{1}^{3} 2\pi y \sqrt{ y^4+2y^2+1}~dy \\\Large \int_{1}^{3} 2\pi y \sqrt{ (y^2)^2+2(y^2)+1}~dy \\ \Large \int_{1}^{3} 2\pi y \sqrt{ (y^2+1)(y^2+1)}~dy \\ \Large \int_{1}^{3} 2\pi y \sqrt{ (y^2+1)^2}~dy \\ \Large \int_{1}^{3} 2\pi y ~ (y^2+1)~dy \\ \Large 2\pi \int_{1}^{3} (y^3+y)~dy $$

OpenStudy (el_arrow):

wow thats alot of y's

OpenStudy (el_arrow):

1/4 y^4 + 1/2 y^2

OpenStudy (el_arrow):

do you have to factor out the 1/4 and 1/2?

OpenStudy (perl):

you shouldn't

OpenStudy (el_arrow):

okay

OpenStudy (el_arrow):

thank you

OpenStudy (el_arrow):

do you do the same steps if the curve rotates about the y-axis?

OpenStudy (perl):

yes

OpenStudy (el_arrow):

okay thanks

OpenStudy (perl):

$$\\ \Large 2\pi \int_{1}^{3} (y^3+y)~dy= 2\pi (\frac{y^4}{4}+\frac{y^3}{3})\LARGE|_1^3 $$

OpenStudy (el_arrow):

@perl one quick question what did you do to the 2 and the 2 exponent on the fourth line?

OpenStudy (perl):

can you point to it , or draw it

OpenStudy (el_arrow):